Consider the electrochemical cell:
Pt(s)|Hg(l)| Hg2Cl2(s)|HCl(aq, 1.0M)|Hg(g,1 bar)|Pt(s)
a.) write the overall balanced redox reaction.
What is the notation saying up above. Please explain very stuck.
b.) dH* > 0 and dS*>0 for the reaction. Upon increasing the temperature, will the cell potential 1.) increase 2.) decrease 3.) stay the same 4.) There is insufficient information to say.
Please explain b. I don't even understand the question!
Answer #1

[The cell looks kind of odd to me! Check again if the cell is like this:

Pt(s)|Hg(l)| Hg2Cl2(s)|HCl(aq, 1.0M)|H2 (g,1 bar)|Pt(s) ; it will be more correct if it is in the reverse order, only then the cell reaction would be a feasible one.]

a.

Pt(s)|Hg(l)| Hg2Cl2(s)|HCl(aq, 1.0M)|Hg(g,1 bar)|Pt(s)

Oxidation reaction at anode:        2Hg(l) + 2Cl-(aq, 1 M) = Hg2Cl2(s) + 2e

Reduction reaction at cathode:   Hg2Cl2(s) + 2e = 2Hg (g) + 2Cl-(aq, 1 M)

____________________________________________________________

Overall reaction :                            2Hg(l) ---> 2Hg (g)             [omit the common terms from both side.]

Note: A single vertical line indicates a phase boundary and two vertical lines indicate that the solutions are connected by a salt bridge eliminating the liquid junction potential. A dotted line shows a porous barrier(with transference) Cells with transference generates a liquid junction potential which is needed to remove by a salt bridge.

Here the cell is a 'cell without transference'.

b. Relation between free energy (dG) and EMF (E) of the cell is: dG = -nFE; where F = Faraday's constant.

and relation between dG, dH and dS is:     dG = dH - TdS

dH and dS is both positive. It implies that TdS also will be positive as temperature cannot be negative. If the temperature is increased value of TdS will increase. So (dH - TdS) i.e., dG will decrease with increasing temperature.

As dG = -nFE, so if dG is decreased, value of E will increase with increasing temperature.